What is bed turnover in relation to hospital data? Furthermore, how does it relate to patient census and occupancy rates? Provide an example.
Answer #1

Bed turnover is a hospital utilization measurements that a number of times changes in occupation of bed during a time period. that is a number of patient discharge and death during a particular time period and average bed count in the period of time.
example:
Number of discharges in the year 2000 is 2100
Number of available bed in 2000 is 500
hospital bed turn over rate is = 2100/500 = 4.2
It has relation to patient census and occupancy rate
Patient census is a number of the patient admitted to the hospital at the given time period per day.
An occupancy rate is a number of days the patient admitted in the hospital in the given time period
Example:
Total number of inpatient census for one week in August is 1510
the average daily census is =1510/7 = 216